# Structural Anisotropy Stabilises Asymmetric Beating in Instability Driven Filaments

### v1

@@ 2.2 Bifurcation and Stability Analyses
The filament model offers a method to generate numerical solutions given an initial configuration. In order to understand the bifurcations between different states that arise from the model, we need to first obtain these states and then assess their stability. We use a Jacobian Free Newton-Kyrlov subspace method (JFNK) to find the stable or unstable steady-state and time-periodic solutions, as described in [50, 51, 37]. JFNK seeks solutions which minimise the error between the initial state and the state after being advanced for a period of time. Therefore, we can only use this to identify steady and time periodic solutions - not, for instance, quasiperiodic or chaotic dynamics. The system is solved using Newton’s method, where the Generalised Residual Method (GMRES) is employed in each Newton iteration to solve for the update. Once we have a solution for a single set of parameters, we can then use continuation to track the solution branch for different values of $f,\kappa_{0}$ or $\beta$, including values of these parameters for which the solution may be unstable.
After finding the steady-state or time-periodic base state, we can analyse its stability numerically. This is achieved by using the Arnoldi method to establish the eigenvalues of the system’s Jacobian, for steady state solutions, or the Floquet matrix exponent, if the state is time-periodic. Further details of this method can be found in [37].
It may be that two solutions are stable for the same values of parameters,. If this is the case, we expect for there to be an unstable state which lies on the boundary of the two stable solutions in the state space. We can identify the unstable solution by using the bisection algorithm outlined in [52], which iteratively bisects between initial conditions leading to either stable state. While this method allows us to identify an unstable solution lying on the boundary surface (separatrix) between the stable states in the state space, we only observe this transiently, before the filament converges to one of the stable states. Furthermore, if we expect multiple unstable solutions to exist for a set of parameters, bisection can only compute a solution stable (or attractive) along the directions tangential to the boundary surface between the two stable states.
@@ 3.1 Planar Dynamics
We begin by considering an isotropic filament with finite intrinsic curvature, i.e. $\beta=1,$ and $\kappa_{0}\neq 0$, with its motion restricted to the plane in which it is deformed by its intrinsic curvature. The corresponding phase diagram of the planar filament states in $\kappa_{0}-f$ space, where the plane is aligned with the filament’s intrinsic curvature, is summarised in Figure 2a. These IVPs reveal that we observe two types of behaviour; steady states below a critical value of the forcing, or self-sustained beating oscillations when this critical value is exceeded, matching observations in [31].
@@
When the force exceeds a critical value, the filament buckles and we observe planar oscillations. To find this critical value for each $\kappa_{0},$ we track the steady state using JFNK and use stability analysis, which reveals that the buckling occurs at $f=f^{*}\approx 35.3,$ independently of $\kappa_{0}$, through a Hopf bifurcation. For $\kappa_{0}=0,$ numerical simulations after the bifurcation show that the planar beating oscillations are symmetric. However, this symmetry is broken by intrinsic curvature; for $\kappa_{0}>0,$ simulations reveal asymmetric beating patterns as shown in Figure 2d, and observed in [31]. Close to the bifurcation this asymmetry is strongest, but further away from the bifurcation, where the follower force is strong, the effect of intrinsic curvature is less pronounced and the beating becomes more symmetric.
Figure 2: (a) Results from 2D numerical simulations of a filament with intrinsic curvature $(\kappa_{0}\geq 0,$ $\beta=1)$. For $f<f^{*}\approx 35.3$, filaments remain in their steady non-trivial state which varies with (b) $\kappa_{0}$ (shown for $f=0,$ $\kappa_{0}\in\{0,\pi/4,\pi/2,\pi/4\}$) and (c) $f$ (shown for $\kappa_{0}=3\pi/4,$ $f\in\{0,5,10,20\}$). For $f>f^{*}$ filaments undergo asymmetric oscillations for $\kappa_{0}>0$, as shown for (d) $\kappa_{0}=3\pi/4$.
@@ 3.2 Non-planar Dynamics - Initial Value Problems
For $\kappa_{0}=0$ [37], we observe the symmetric whirling state immediately after the buckling; a rigid body rotation of the centerline, with the filament tip tracing a circle over one period. Increasing the forcing further, whirling is replaced by a quasiperiodic solution which we term QP1 - an elliptical whirling which rotates unidirectionally. QP1 only exists for a small region of $f,$ before being replaced with planar beating. Planar beating is also observed in the 2D simulations, but is now symmetric (as $\kappa_{0}=0$) and can occur in any plane (due to the rotational symmetry of the problem).
Increasing $\kappa_{0}>0,$ the initial buckling occurs at smaller values of the forcing. The whirling behaviour still exists but is no longer symmetric, as shown in Figure 3b and Supplemental Video 1. In fact, the asymmetry is strongest as we increase the intrinsic curvature, and choose forcing values closer to the initial bifurcation. Increasing the forcing has the effect of smoothing out the asymmetry, as the follower force dominates over the torques maintaining the intrinsic curvature.
Increasing $f$ further, we observe that whirling gives way to a new periodic solution for $\kappa_{0}>0$, which we call P1. This consists of a bent-over beating, where the filament is bent in the plane provided by the intrinsic curvature and beats lies in the plane perpendicular to it, as shown in Figure 3a inset and Supplemental Video 2. In Figure 3c, we fix $f=200$ and increase $\kappa_{0}$ from 00 to $43\pi/$ 4 to show how P1 continues from the planar beating solution. We see that increasing $\kappa_{0}$ increases the maximum perpendicular distance of the tip from the beat plane. Varying $\kappa_{0}>0$, planar beating is no longer visible with IVPs alone - it appears that the 2D asymmetric beating patterns observed in the planar case are unstable in the 3D model.
@@
Figure 3: (a) Results of numerical simulations of a filament with intrinsic curvature in 3D $(\kappa_{0}\geq 0,\beta=1)$. (b) Snapshots of filaments over one period undergoing whirling dynamics for two values of intrinsic curvature close to ($f=40$) and further from ($f=100$) the bifurcation. We see asymmetries in whirling are more pronounced close to the bifurcation, and for higher $\kappa_{0}.$ (c) Snapshots of filaments over one period for $f=200,$ increasing $\kappa_{0}$ from 00 to $43\pi/$ 4. The dynamics change from planar beating for $\kappa_{0}=0$ (left) to P1 for $\kappa_{0}>0$ (right).
@@ 3.3 Bifurcation and Stability Analyses
Figure 4: Bifurcation diagrams for an isotropic filament ($\beta=1$) with intrinsic curvature for (a) $\kappa_{0}=0$, (b) $\kappa_{0}=\pi/4$, (c) $\kappa_{0}=\pi/2$ and (d) $\kappa_{0}=3\pi/4$. Dotted/full lines refer to unstable/stable solutions. Inset, left, shows the bifurcation diagram near the initial buckling event and inset, right, shows the bifurcation diagram around the bistable regions, if these are present.
Figure 5: The real part of the four largest eigenvalues associated to the steady state for $\beta=1$, (a) $\kappa_{0}=0$, and (b) $\kappa_{0}=3\pi/4$. Analysing the corresponding eigenvectors, shown on the right, allows us to identify the branches corresponding to off-plane and in-plane beating. (c) The real part of the largest eigenvalue associated to the steady state for various values of $\kappa_{0}.$ We see that the eigenvalue becomes unstable for smaller $f$ as we increase $\kappa_{0}.$ (d) The real part of the largest eigenvalue associated to the P1 state for various values of $\kappa_{0}$. For $\kappa_{0}>\kappa_{0}^{*}$, we see the eigenvalue remains stable.
To explore in more detail how filament dynamics change with $\kappa_{0},$ we track the steady and time-periodic states with JFNK and perform a stability analysis to establish bifurcations within the system. An overview of the resulting bifurcation diagrams for $f\in[0,200]$ is shown in Figure 4. We now discuss the key changes in the bifurcation diagrams as we increase $\kappa_{0}$. We provide a more detailed version of this discussion in Appendix B.1.
@@ 3.3.1 Isotropic Case
We begin by recapping the isotropic case, where $\kappa_{0}=0$, shown in Figure 4a. Focusing first on the initial buckling event, in [37, 38] it was shown that two complex conjugate pairs of eigenmodes become unstable at $f=f^{*}$ (see Figure 5a). We hence identify this bifurcation as a Double Hopf bifurcation. The first complex conjugate pair is associated to beating in the $(x,z)-$plane, and the second is associated to beating in the perpendicular plane, as shown by the eigenvectors in Figure 5a (right). It is these modes that combine to give rise to the stable whirling solution observed after the bifurcation, as well as the unstable planar beating branch. For larger $f$, whirling becomes unstable and gives rise to QP1. The QP1 branch exists for a small region of $f$, and vanishes when planar beating stabilises for larger $f$ (see Figure 4a, inset). Further discussion on the bifurcation diagram in the isotropic case, including exploration of higher forcing values, can be found in [37].
@@ 3.3.2 Initial Buckling Event
As we increase $\kappa_{0}$, we see a change in the initial buckling event; as shown in Figure 4a-d left inset, the initial buckling happens at smaller values of $f$ as we increase $\kappa_{0}.$ We also now observe two solution branches bifurcating from the steady state at different values of the forcing, and note that whirling no longer bifurcates from the steady state for $\kappa_{0}>0$. Instead, we find the P1 branch bifurcates from the steady state initially at a critical forcing $f<f^{*}$, followed by the unstable planar beating branch at $f=f^{*}$. To explore these changes in the bifurcation diagrams, we analyse the eigenmodes associated to the steady state.
We plot the real parts of the four dominant eigenmodes of the steady state for $\kappa_{0}=3\pi/4$ in Figure 5b. We see that the two modes corresponding to buckling in either plane separate; the mode corresponding to in-plane, asymmetric beating still becomes unstable at $f=f^{*},$ whereas the second mode, corresponding to off-plane beating, becomes unstable earlier. Hence for $\kappa_{0}>0$, at the initial bifurcation, it is only the off-plane modes that are becoming unstable which lead to P1 through a Hopf bifurcation. In Figure 5c, we plot the real part of the dominant eigenmode for various values of $\kappa_{0}$. We see that the off-plane modes become unstable earlier as we increase the intrinsic curvature, meaning the steady state buckles at smaller forcing values, as we observed earlier by solving IVPs.
@@
Immediately following the initial buckling bifurcation for $\kappa_{0}>0$, we see from Figures 4b-d that the P1 solution is stable. For moderate $\kappa_{0}$, the P1 solution becomes unstable for larger $f$, giving way to the stable whirling solution (see Figures 4(b-c), left inset). This can again be linked to the initial buckling event. As the whirling solution is known to arise from a combination of both the in-plane and off-plane buckling eigenmodes [37], it follows that the whirling solution is always born after the second buckling mode becomes unstable, branching from the P1 solution branch.

### v2

@@ 2.2 Bifurcation and Stability Analyses
The filament model offers a method to generate numerical solutions given an initial configuration. In order to understand the bifurcations between different states that arise from the model, we need to first obtain these states and then assess their stability. We use a Jacobian Free Newton-Kyrlov subspace method (JFNK) to find the stable or unstable steady-state and time-periodic solutions, as described in (Viswanath2007RecurrentTurbulence, ; Willis2019EquilibriaThem, ; Clarke2024BifurcationsFilaments, ). JFNK seeks solutions which minimise the error between the initial state and the state after being advanced for a period of time. Therefore, we can only use this to identify steady and time periodic solutions - not, for instance, quasiperiodic or chaotic dynamics. The system is solved using Newton’s method, where the Generalised Residual Method (GMRES) is employed in each Newton iteration to solve for the update. Once we have a solution for a single set of parameters, we can then use continuation to track the solution branch for different values of $f,\kappa_{0}$ or $\beta$, including values of these parameters for which the solution may be unstable.
After finding the steady-state or time-periodic base state, we can analyse its stability numerically. This is achieved by using the Arnoldi method to establish the eigenvalues of the system’s Jacobian, for steady state solutions, or the Floquet matrix exponent, if the state is time-periodic. Further details of this method can be found in (Clarke2024BifurcationsFilaments, ).
It may be that two solutions are stable for the same values of parameters. If this is the case, we expect for there to be an unstable state which lies on the boundary of the two stable solutions in the state space. We can identify the unstable solution by using the bisection algorithm outlined in (Skufca2006EdgeFlow, ), which iteratively bisects between initial conditions leading to either stable state. While this method allows us to identify an unstable solution lying on the boundary surface (separatrix) between the stable states in the state space, we only observe this transiently, before the filament converges to one of the stable states. Furthermore, if we expect multiple unstable solutions to exist for a set of parameters, bisection can only compute a solution stable (or attractive) along the directions tangential to the boundary surface between the two stable states.
@@ 3.1 Planar Dynamics
Figure 2: (a) Results from 2D numerical simulations of a filament with intrinsic curvature $(\kappa_{0}\geq 0,$ $\beta=1)$. The black line is drawn to indicate the approximate boundaries of the solution regions. For $f<f^{*}\approx 35.3$, filaments remain in their steady non-trivial state which varies with (b) $\kappa_{0}$ (shown for $f=0,$ $\kappa_{0}\in\{0,\pi/4,\pi/2,\pi/4\}$) and (c) $f$ (shown for $\kappa_{0}=3\pi/4,$ $f\in\{0,5,10,20\}$). For $f>f^{*}$ filaments undergo asymmetric oscillations for $\kappa_{0}>0$, as shown for (d) $\kappa_{0}=3\pi/4$.
We begin by considering an isotropic filament with finite intrinsic curvature, i.e. $\beta=1,$ and $\kappa_{0}\neq 0$, with its motion restricted to the plane in which it is deformed by its intrinsic curvature. The corresponding phase diagram of the planar filament states in $\kappa_{0}-f$ space, where the plane is aligned with the filament’s intrinsic curvature, is obtained by running one initial value problem for each set of parameters. Each initial value problem has an initial condition of a vertically upright filament with random, small perturbations to the forcing on a subset of the segments. We summarise the results in Figure 2a. For this study we consider $\kappa_{0}\in[0,3\pi/4]$. This range of values is appropriate as it includes values used in the literature in numerical studies (Wang2023Generalized-NewtonianFilament, ), in particular to recover dynamics similar to those observed experimentally (Chakrabarti2019HydrodynamicFilaments, ; Sartori2016DynamicFlagella, ). These IVPs reveal that we observe two types of behaviour; steady states below a critical value of the forcing, or self-sustained beating oscillations when this critical value is exceeded, matching observations in (Wang2023Generalized-NewtonianFilament, ).
@@
When the force exceeds a critical value, the filament buckles and we observe planar oscillations. To find this critical value for each $\kappa_{0},$ we track the steady state using JFNK and use stability analysis. For $\kappa_{0}=0,$ this reveals that the buckling occurs at $f=f^{*}\approx 35.3,$ through a Hopf bifurcation. For $\kappa_{0}>0$, the bifurcation occurs through the same mechanism, but we observe a small increase in the critical forcing of approximately $0.5\%$ from $\kappa_{0}=0$ to $\kappa_{0}=3\pi/4.$ For $\kappa_{0}=0,$ numerical simulations after the bifurcation show that the planar beating oscillations are symmetric. However, this symmetry is broken by intrinsic curvature; for $\kappa_{0}>0,$ simulations reveal asymmetric beating patterns as shown in Figure 2d, and observed in (Wang2023Generalized-NewtonianFilament, ). We plot the frequencies of these solutions for various $\kappa_{0}$ in Appendix C. Close to the bifurcation this asymmetry is strongest, but further away from the bifurcation, where the follower force is strong, the effect of intrinsic curvature is less pronounced and the beating becomes more symmetric.
@@ 3.2 Non-planar Dynamics - Initial Value Problems
Figure 3: (a) Results of numerical simulations of a filament with intrinsic curvature in 3D $(\kappa_{0}\geq 0,\beta=1)$. The black lines are drawn to indicate the approximate boundaries of the solution regions. (b) Snapshots of filaments over one period undergoing whirling dynamics for three values of intrinsic curvature close to ($f=40$) and further from ($f=100$) the bifurcation. We see asymmetries in whirling for $\kappa_{0}>0$ are more pronounced close to the bifurcation, and for higher $\kappa_{0}.$ (c) Snapshots of filaments over one period for $f=200,$ increasing $\kappa_{0}$ from 0 to $3\pi/4$. The dynamics change from planar beating for $\kappa_{0}=0$ (left) to P1 for $\kappa_{0}>0$ (right).
@@
For $\kappa_{0}=0$ (Ling2018Instability-drivenMicrofilaments, ; Clarke2024BifurcationsFilaments, ), we observe the symmetric whirling state immediately after the buckling; a rigid body rotation of the centerline, with the filament tip tracing a circle over one period. Increasing the forcing further, whirling is replaced by a quasiperiodic solution which we term QP1 - an elliptical whirling which rotates unidirectionally. QP1 only exists for a small region of $f,$ before being replaced with planar beating. Planar beating is also observed in the 2D simulations, but is now symmetric (as $\kappa_{0}=0$) and can occur in any plane (due to the rotational symmetry of the problem).
Increasing $\kappa_{0}>0,$ the initial buckling occurs at smaller values of the forcing. The whirling behaviour still exists but is no longer symmetric, as shown in Figure 3b and Supplemental Video 1. The frequencies of the whirling solutions are plotted in Appendix C. In fact, the asymmetry is strongest as we increase the intrinsic curvature, and choose forcing values closer to the initial bifurcation. Increasing the forcing has the effect of smoothing out the asymmetry, as the follower force dominates over the torques maintaining the intrinsic curvature.
Increasing $f$ further, we observe that whirling gives way to a new periodic solution for $\kappa_{0}>0$, which we call P1. This consists of a bent-over beating, where the filament bends in the plane it would deform in the absence of the follower force, and beats in the plane perpendicular to it, as shown in Figure 3a inset and Supplemental Video 2. The intrinsic curvature breaks left-right symmetry, in such a way that the filament beats in the left half of the plane. In Figure 3c, we fix $f=200$ and increase $\kappa_{0}$ from 0 to $3\pi/4$ to show how P1 continues from the planar beating solution. We see that increasing $\kappa_{0}$ increases the maximum perpendicular distance of the tip from the beat plane. We plot the frequencies of P1 for various $f$ and $\kappa_{0}$ in Appendix C. Varying $\kappa_{0}>0$, planar beating is no longer visible with IVPs alone - it appears that the 2D asymmetric beating patterns observed in the planar case are unstable in the 3D model.
@@ 3.3 Bifurcation and Stability Analyses
Figure 4: Bifurcation diagrams showing the different solutions as we vary $f$, and the corresponding mean bending energy of these solutions, $\bar{E}_{b}$, for an isotropic filament ($\beta=1$) with intrinsic curvature for (a) $\kappa_{0}=0$, (b) $\kappa_{0}=\pi/4$, (c) $\kappa_{0}=\pi/2$ and (d) $\kappa_{0}=3\pi/4$. Dashed/full lines refer to unstable/stable solutions. Square markers are used to indicate the location of bifurcations as we vary $f$. Inset, left, shows the bifurcation diagram near the initial buckling event and inset, right, shows the bifurcation diagram around the bistable regions, if these are present. As $\kappa_{0}$ increases we see the emergence of the P1 solution branch (b-d), from which whirling bifurcates (b and c, left inset). We observe bistability between P1 and whirling (b and c, right inset) and, for the largest values of $\kappa_{0}$, whirling vanishes completely (d).
Figure 5: The real part of the four largest eigenvalues, $\lambda_{i}$ for i=1,,4i=1,...,4, associated to the steady state for $\beta=1$, (a) $\kappa_{0}=0$, and (b) $\kappa_{0}=3\pi/4$. Analysing the corresponding eigenvectors at each bifurcation, obtained as described in the main text and displayed on the right, allows us to identify the branches corresponding to off-plane and in-plane beating. (c) The real part of the largest eigenvalue associated to the steady state for various values of $\kappa_{0}.$ We see that the eigenvalue becomes unstable for smaller $f$ as we increase $\kappa_{0}.$ (d) The real part of the largest eigenvalue associated to the P1 state for various values of $\kappa_{0}$. For $\kappa_{0}>\kappa_{0}^{*}$, we see the eigenvalue remains stable.
To explore in more detail how filament dynamics change with $\kappa_{0},$ we track the steady and time-periodic states with JFNK and perform a stability analysis to establish bifurcations within the system. An overview of the resulting bifurcation diagrams for $f\in[0,200]$ is shown in Figure 4, where we plot the non-dimensionalised mean bending energy, given by
$\bar{E}_{b}=\frac{1}{T}\int_{0}^{T}\int_{0}^{L}\frac{1}{2}\kappa(s)^{2}\;ds\;dt,$ (13)
for each solution branch for various values of $\kappa_{0}$. Here $\kappa(s)$ is the local curvature of the filament. While the bending energy is usually defined with respect to the intrinsic curvature, we instead use the standard bending energy described above for ease of comparison as we vary $\kappa_{0}$. We now discuss the key changes in the bifurcation diagrams as we increase $\kappa_{0}$. We provide a more detailed version of this discussion in Appendix B.1.
@@ 3.3.1 Isotropic Case
We begin by recapping the isotropic case, where $\kappa_{0}=0$, shown in Figure 4a. Focusing first on the initial buckling event, in (Clarke2024BifurcationsFilaments, ; Schnitzer2025OnsetFilament, ) it was shown that two complex conjugate pairs of eigenmodes become unstable at $f=f^{*}$ (see Figure 5a). We hence identify this bifurcation as a Double Hopf bifurcation.
To understand the buckling bifurcation further, we visualise the eigenvectors. Defining each eigenvalue-eigenvector pair as $(\lambda_{i},\bm{x}_{i})$, we note that the $\bm{x}_{i}$ are complex for i=1,,4i=1,...,4. Therefore, to interpret their physical meaning, we take linear combinations of the eigenvectors, $\bm{v}_{1}=a(\bm{x}_{1}+\bm{x}_{2}),\bm{v}_{2}=a(\bm{x_{1}}-\bm{x}_{2})$ and similar for $\bm{v}_{3}$ and $\bm{v}_{4}$ using $\bm{x}_{3}$ and $\bm{x}_{4}$, where $a$ is an arbitrary constant chosen to emphasise the structure of the eigenmode. As the magnitude of the state variable relates to angles of rotation of the local basis, taking larger values of $a$ results in more exaggerated filament deformation. We show these eigenvectors in Figure 5 for $a=5.$ We observe that the first complex conjugate pair is associated to beating in the $(x,z)$-plane, and the second is associated to beating in the perpendicular plane, as shown by the eigenvectors in Figure 5a (right). It is these modes that combine to give rise to the stable whirling solution observed after the bifurcation, as well as the unstable planar beating branch. For larger $f$, whirling becomes unstable and gives rise to QP1. The QP1 branch exists for a small region of $f$, and vanishes when planar beating stabilises for larger $f$ (see Figure 4a, inset). Further discussion on the bifurcation diagram in the isotropic case, including exploration of higher forcing values, can be found in (Clarke2024BifurcationsFilaments, ).
@@ 3.3.2 Initial Buckling Event
As we increase $\kappa_{0}$, we see a change in the initial buckling event; as shown in Figure 4a-d left inset, the initial buckling happens at smaller values of $f$ as we increase $\kappa_{0}.$ We also now observe two solution branches bifurcating from the steady state at different values of the forcing, and note that whirling no longer bifurcates from the steady state for $\kappa_{0}>0$. Instead, we find the P1 branch bifurcates from the steady state initially at a critical forcing $f<f^{*}$, followed by the unstable planar beating branch at $f\approx f^{*}$. To explore these changes in the bifurcation diagrams, we analyse the eigenmodes associated to the steady state.
We plot the real parts of the four dominant eigenmodes of the steady state for $\kappa_{0}=3\pi/4$ in Figure 5b, including visualisations of the eigenvectors obtained as discussed previously. We see that the two modes corresponding to buckling in either plane separate; the mode corresponding to in-plane, asymmetric beating still becomes unstable at $f\approx f^{*},$ whereas the second mode, corresponding to off-plane beating, becomes unstable earlier. We note that the off-plane eigenmodes do not both lie in the $(y,z)$-plane. This is because the intrinsic curvature breaks the symmetry about the $(y,z)$-plane, and so we expect one of these eigenmodes to reflect this broken symmetry. This can be seen in the off-plane eigenvectors (see Figure 5b, top right). Hence for $\kappa_{0}>0$, at the initial bifurcation, it is only the off-plane modes that are becoming unstable which lead to P1 through a Hopf bifurcation. In Figure 5c, we plot the real part of the dominant eigenmode for various values of $\kappa_{0}$. We see that the off-plane modes become unstable earlier as we increase the intrinsic curvature, meaning the steady state buckles at smaller forcing values, as we observed earlier by solving IVPs.
@@
Immediately following the initial buckling bifurcation for $\kappa_{0}>0$, we see from Figures 4b-d that the P1 solution is stable. For moderate $\kappa_{0}$, the P1 solution becomes unstable for larger $f$, giving way to the stable whirling solution (see Figures 4(b-c), left inset). This can again be linked to the initial buckling event. As the whirling solution is known to arise from a combination of both the in-plane and off-plane buckling eigenmodes (Clarke2024BifurcationsFilaments, ), it follows that the whirling solution is always born after the second buckling mode becomes unstable, branching from the P1 solution branch.